OpenStudy (anonymous):

lim x -> 0 sin5x / 2x( 1+ 4 cos 2x ) . help !

OpenStudy (loser66):

\[\lim_{x \rightarrow 0}\frac{\sin5x*5}{5*2x(1+4\cos2x)}\]

OpenStudy (loser66):

and split down to (let lim sign aside) = \[\frac{sin5x}{5} * \frac{5}{2x(1+4cos2x)}\]

OpenStudy (loser66):

and lim of the first term =1 so, you just have limit of the second one =infinitive

OpenStudy (anonymous):

hah ? why math error ?

OpenStudy (loser66):

dont get

OpenStudy (anonymous):

but i got 1/2 . i just want to make sure mine is right .

OpenStudy (loser66):

how to get 1/2?

OpenStudy (loser66):

you have the form 0/0, you must manipulate to avoid it.

OpenStudy (loser66):

or you can use l'Hopital rule. Either way is ok

OpenStudy (anonymous):

huh ? never heard of it

OpenStudy (loser66):

ok, show me your work,

OpenStudy (anonymous):

wait up

OpenStudy (anonymous):

@Loser66 your process even ends up with 1/2. :)

OpenStudy (anonymous):

sin5x /5x . 5x / 2x( 1 + 4 cos 2x ) . cut the x's at 5/2 . because sin5x/5x = 1

OpenStudy (loser66):

ah yes, I got it. thanks for pointing out my mistake,

OpenStudy (anonymous):

then 1 x (5/ (2x0 ) ( 1 + 4 cos 2(0 )) = 1/2

OpenStudy (loser66):

@bellehunny I am sorry.!! I did it intuitively,

OpenStudy (anonymous):

ohhh , no wonder

OpenStudy (loser66):

I didn't look up at my note , just remember something like that.... and made a dumb mistake

OpenStudy (anonymous):

it's okay , everybody makes mistake . hehe

OpenStudy (loser66):

good, quite fair

OpenStudy (anonymous):

btw , did i do it right ?